If you have 45.6 g NH3, how many grams of F2 are required for a complete reaction? Need to show ALL calculation work.


1
Expert's answer
2021-02-01T03:46:40-0500

he balanced reaction equation is:

2NH3 + 3F2 "\rightarrow" N2 + 6HF.


In this reaction, 2 moles of ammonia react with 3 moles of fluorine:

"\frac{n(NH_3)}{2} = \frac{n(F_2)}{3}" .


The number of the moles of ammonia is its mass "m" divided by its molar mass "M" =17.03 g/mol:

"n(NH_3) = \frac{45.6}{17.03} = 2.678" mol.


Then, the number of the moles of F2 needed for the complete reaction is:

"n(F_2) = 3\frac{n(NH_3)}{2} = 4.016" mol.


Finally, the mass of the fluorine is its number of the moles times its molar mass 38.00 g/mol:

"m(F_2) = n\u00b7M = 4.016\u00b738.00 = 152.62" g.


Answer: if you have 45.6 g NH3, 152.62 grams of F2 are required for a complete reaction.